# What makes math problems hard for reinforcement learning: a case study

### v1

@@ 3.3. Comparison
could be related neither to $\text{AK}(3)$ nor to the trivial presentation with any sequence of moves that allowed the length of each relator to increase up to 20.
@@
We also plot the lengths of AC sequences discovered by greedy search as functions of $n$ and the maximum increase in the presentation length (Figure 3). Unsurprisingly, path lengths increase proportionally with the increase in the length of the presentation (3(a)). The following presentation with $n=5$ had the longest AC trivialization path,
$\langle x^{-1}y^{5}x=y^{6}\mid x=yx^{2}y^{-1}\rangle$
requiring a sequence of 344 AC-moves. Note that greedy search does not necessarily find the shortest paths of trivialization. We will see in Subsection 4.3 that a Reinforcement Learning algorithm finds shorter trivializing sequences for many examples of the Miller–Schupp series. This again hints at the potential utility of exploring more efficient criteria for ordering nodes in the priority queue.
@@
3.4. Limitations
@@
3.5. Proof of Theorem 1
As mentioned in Section 1, one nice byproduct of our analysis is that the shortest mysterious AC presentation, namely $\text{AK}(3)$, is stably AC-trivial. The goal of this part is to present a proof of this statement.
First, in order to make this part of the paper self-contained, let us remind the reader that the term “stable” (a.k.a. “weak”) refers to one of many variants of the Andrews–Curtis conjecture, see e.g. [MMS, Meier2016, Bagherifard2021], where in addition to the usual AC-moves one is allowed to use two more transformations:
  1. (AC4)
  2. (AC5)
Definition 2.
If two balanced presentations of the trivial group are related by a sequence of AC-transformations (AC1) through (AC5), we say that they are stably AC-equivalent.
The stable Andrews–Curtis conjecture states that any balanced presentation is stably AC-equivalent to the trivial presentation. To the best of our knowledge, prior to this work, the shortest potential counterexample to the standard Andrews–Curtis conjecture, $\text{AK}(3)$, was also a potential counterexample to the stable Andrews–Curtis conjecture. Our proof that $\text{AK}(3)$ is stably AC-trivial builds on the following result.
Theorem (Myasnikov, Myasnikov, and Shpilrain, [MMS]).
Using the notation $[a,b]=aba^{-1}b^{-1}$ and $[a,b,c]=[[a,b],c]$, any presentation of the following form is a presentation of the trivial group:
$\langle x,y,z\mid x=z\cdot[y^{-1},x^{-1},z],\ y=x\cdot[y^{-1},x^{-1},z^{-1}]\cdot[z^{-1},x],\ w\rangle,$
where $w$ is a word in $x$, $y$, and $z$ whose exponent sum on $x$, $y$, and $z$ equals ±$\pm 1$. Moreover, all such presentations are stably AC-trivial.
These presentations are obtained by applying Reidemeister moves to the knot diagram of the unknot and using the fact that Reidemeister moves applied to a knot diagram give stably AC-equivalent Wirtinger presentations of the knot group, cf. [WADA1994241].
For $w=x^{-1}yz$, one of the relators eliminates the generator $z$, resulting in the following length 25 presentation with two generators:
$\langle x,y\mid x^{-1}y^{-1}xy^{-1}x^{-1}yxy^{-2}xyx^{-1}y,\ y^{-1}x^{-1}y^{2}x^{-1}y^{-1}xyxy^{-2}x\rangle.$
We discovered a sequence of AC-transformations (AC1)–(AC5) that relates this presentation to $\text{AK}(3)$. This also makes $\text{AK}(3)$ the shortest stably AC-trivial presentation that is not yet known to be AC-trivial. It is plausible that by varying $w$ one can show that other presentations of the Akbulut–Kirby series (or the Miller–Schupp series) are also stably AC-trivial. We leave this question for future work.
Specifically, using search algorithms described earlier in this section we placed a cutoff of a maximum of 1 million nodes to visit for each of our search algorithms and allowed the length of each relator to increase up to $15$. The greedy search found a path connecting this presentation to $\text{AK}(3)$, while breadth-first search could only reduce the presentation’s length to $14$. We repeated the search process with breadth-first search with a cutoff of 5 million nodes. It failed to reduce the presentation length any further.
The sequence of moves that connects the length-25 presentation to $\text{AK}(3)$ can be conveniently expressed in terms of the following 12 transformations:
among which a careful reader can recognize moves (AC1) and (AC2) introduced in Section 2. Expressed in terms of the moves $h_{i}$, the desired sequence has length 53 and looks as follows:
This sequence should be read from left to right; first apply $h_{9}$, then $h_{7}$, and so forth. This follows the standard convention of many programming languages, which iterate over lists from left to right by default. The length of the presentation did not exceed $25$ during its path to $\text{AK}(3)$. We do not know if this is the shortest path between the two presentations.
@@ 4. Reinforcement Learning
This section is organized as follows: in Subsection 4.1, we discuss how the problem underlying the Andrews–Curtis conjecture can be formulated as a Markov Decision Process. In Subsection 4.2, we discuss details of a specific reinforcement learning algorithm, called the Proximal Policy Optimization algorithm, which we used to find AC trivializations of balanced presentations. Finally, in Subsection 4.3, we discuss the results of our work, comparing the performance of PPO with that of the classical search algorithms studied in the previous section.
@@ 4.1. Markov Decision Process
The schematic picture of how these objects interact with each other is as follows. We start with a state $s_{0}$ sampled from the distribution $\rho$ and take an action $a_{0}$. This results in a state $s_{1}$ with probability $P(s_{1}\mid s_{0},a_{0})$. The transition gets a “reward” $r_{0}=R(s_{0},a_{0},s_{1})$ which quantifies the effectiveness of the action in contributing toward achieving an ultimate goal. From state $s_{1}$, we repeat this process, obtaining a trajectory of states
@@
For a given problem at hand, we may not a priori know the actions $\{a_{t}\}$ and states $\{s_{t+1}\}$ that maximize the return. Deep reinforcement learning presents a solution to this problem: we train a neural network that learns a map from states to actions with the objective of maximizing the cumulative return. More precisely, we learn a map called the “policy” function $\pi\colon S\to\mathcal{P}(A)$ that assigns to each state a probability distribution over actions. At time step $t$, an action $a_{t}\sim\pi(\cdot\mid s_{t})$ is sampled that gives the next state $s_{t+1}$.555 In general, we need to specify the probability transition function $P(\cdot\mid s_{t},a_{t})$ from which the next state would be sampled. In this paper, we take this probability distribution to be a delta function centered at a single state, which we write as $a_{t}(s_{t})$. In the next subsection we discuss the specific algorithm and the objective function that we used in our study.
@@ 4.2. Proximal Policy Optimization
It turns out that this update depends on the gradient of the logarithm of the policy function itself and an “advantage function” $A^{\pi}(s,a)$ which quantifies the relative benefit of taking action $a$ in state $s$ under policy $\pi$ [SpinningUp2018].666 Mathematically, the advantage function is the difference between “on-policy action-value function” $Q^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)\mid s_{0}=s,a_{0}=a]$ and the “on-policy value function”, $V^{\pi}(s)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)\mid s_{0}=s]$. These functions give an estimate of the cumulative return given that the current state-action pair is $(s,a)$ in the case of $Q^{\pi}$ and the current state is $s$ in the case of $V^{\pi}$. Explicitly,
@@
Here $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\text{old}}(a_{t}|s_{t})}$ represents the probability ratio, and $\epsilon$ is a small positive constant (commonly set around 0.1 or 0.2) that controls the clipping value. The raw and the clipped ratio ensure that excessively large policy updates are curtailed, making the optimization process more stable.
The advantage function is estimated during the optimization process using the Generalized Advantage Estimation method of [schulman2018highdimensional]. This method requires an estimation of the on-policy value function (c.f. footnote 6). In PPO, the actor-critic framework is used for this purpose. We train two neural networks in parallel, where the “actor” learns the policy $\pi_{\theta}$ and the “critic” learns the value function.777 Sometimes, these neural networks are made to share their parameters, which helps in stability during training. We experimented with shared as well as unshared parameters and observed better performance by keeping the parameters independent. The value loss $L^{V}$ is the mean squared error between the values for a state as estimated by the critic network before and after a gradient update. Lastly, PPO adds entropy $S$ of the action-space distribution to the full loss function to prevent premature convergence to a suboptimal policy. The full loss function is as follows,
@@
In policy gradient algorithms such as PPO, we alternate between collecting data through the current policy in the “rollout phase” and updating the policy weights through gradient descent in the “optimization phase”. There are various hyperparameters associated to these phases, which we briefly describe now. The rollout phase involves generating multiple trajectories of the underlying Markov Decision Process in parallel using the current policy function. The hyperparameters of this phase include the number of parallel actors $N$, the rollout length $T^{\prime}$, i.e. the number of steps each actor takes, the discount factor $\gamma$ (c.f. Subsection 4.1) and a bootstrapping hyperparameter of Generalized Advantage Estimation $\lambda_{\text{GAE}}$. The dataset of $N\times T^{\prime}$ examples collected in this phase is then used to update the policy. In each epoch, the dataset is shuffled and split into mini-batches before performing gradient descent. In addition to the optimizer hyperparameters (such as learning rate schedule), the number of epochs and the size of mini-batches are hyperparameters of the optimization process. Further details of hyperparameters are summarized in Appendix A.
4.3. Application to the Andrews–Curtis Conjecture
The set of all balanced presentations of the trivial group with a fixed number of generators and the set of AC-transformations play the roles of sets $S$ and $A$, respectively. (Recall notations introduced in Subsection 4.1). Once we choose a reward function $R$ and an initial state distribution $\rho$, we may use the Proximal Policy Optimization algorithm to learn the policy function $\pi$. We tested a few different reward functions in our experiments, observing that the following candidate led to the best performance and stability in training.

### v2

@@ 3.3. Comparison of performance on Miller–Schupp series
could be related neither to $\text{AK}(3)$ nor to the trivial presentation through any path involving only presentations with both relators having length less than 20.
@@
We also plot the lengths of AC sequences discovered by greedy search as functions of $n$ and the maximum increase in the presentation length (Figure 3). Unsurprisingly, path lengths increase proportionally with the increase in the length of the presentation (Figure 3(a)). The following presentation with $n=5$ had the longest AC trivialization path,
$\langle x,y\mid x^{-1}y^{5}x=y^{6},x=yx^{2}y^{-1}\rangle$
requiring a sequence of 344 AC-moves. Note that greedy search does not necessarily find the shortest paths of trivialization. We will see in Section 4.3 that a Reinforcement Learning algorithm finds shorter trivializing sequences for many examples of the Miller–Schupp series. This again hints at the potential utility of exploring more efficient criteria for ordering nodes in the priority queue.
@@
3.4. AC-triviality of $\text{MS}(1,w)$
As mentioned in the previous subsection, the greedy search could solve all $n=1$ presentations of the Miller–Schupp series in our dataset. This motivated us to prove AC-triviality for all $\text{MS}(1,w)$.
In this and the following subsections, we frequently use the substitution transformation mentioned in Section 2.
Theorem 2.
$\text{MS}(1,w)$ is AC-trivial for all $w$.
Proof.
We have
$\text{MS}(1,w)=\langle x,y\mid x^{-1}yx=y^{2},x=w\rangle$
where $w$ has exponent sum 0 in $x$. We can re-write the first relator in the following four ways:
  1. (i)
    $yx=xy^{2}$
  2. (ii)
    $x^{-1}y=y^{2}x^{-1}$
  3. (iii)
    $x^{-1}y^{-1}=y^{-2}x^{-1}$
  4. (iv)
    $y^{-1}x=xy^{-2}$
We can substitute these equations in the second relator to move around all occurrences of $x$ and $x^{-1}$: relations (i) and (iv) move $x$ to the left, relations (ii) and (iii) move $x^{-1}$ to the right. We continue until the second relator becomes $x^{a-1}y^{b}x^{-a}$ for some $a,b\in\mathbb{Z}$. Through conjugation, this simplifies to $x=y^{b}$, which when substituted into the first relator leads to a trivialization of the presentation. ∎
The following result follows from this theorem.
Theorem 3.
$\text{MS}(n,w_{\star})$ for $w_{\star}=y^{-1}xyx^{-1}$ is AC-trivial for all $n$.
Proof.
In the presentation,
$\text{MS}(n,w_{\star})=\langle x,y\mid x^{-1}y^{n}x=y^{n+1},x=y^{-1}xyx^{-1}\rangle$
we can rewrite the second relation as $y^{-1}xy=x^{2}$, and apply the automorphism $x\leftrightarrow y$ to get
$\text{MS}(n,w_{\star})=\langle x,y\mid x^{-1}yx=y^{2},y^{-1}x^{n}y=x^{n+1}\rangle.$
This presentation is the same as $\text{MS}(1,y^{-1}x^{n}yx^{-n})$, which is AC-trivial by Theorem 2. ∎
AC-trivializations of $\text{MS}(n,w_{\star})$ for $n=3,4,5,6,7,8$ were recently obtained using automated theorem proving in [new-ac-for-ms]. Here, we have obtained AC-trivializations of this family for all $n$.
3.5. Length reduction for $\text{AK}(n)$
In Section 3.3, we discussed the performance of classical search algorithms for presentations of the Miller–Schupp series. But what if we apply these algorithms for presentations of the Akbulut–Kirby series instead? We find that starting from $\text{AK}(5)$—a presentation of length $17$—and performing greedy search for $20$ million nodes, reduces its length to 16. Examining the sequence of AC moves connecting the two presentations, we found the following result.
Theorem 4.
For every $n\geq 2$, $\text{AK}(n)$ is AC-equivalent to the presentation
$\langle x,y\mid x^{-1}yx=xyx^{-1}y\ ,\ xy^{n-1}x=yxy\rangle,$
of length $n+11$. This gives reduction in length of $AK(n)$ for all $n\geq 5$.
To prove this theorem, we will use the following result due to [MMS], and prove two additional theorems (Theorem 6 and Theorem 7). From these results, Theorem 4 follows immediately.
Proposition 5 (Myasnikov, Myasnikov, and Shpilrain, [MMS]).
For all $n\geq 2$, $AK(n)$ is AC-equivalent to $\text{MS}(n,w_{1})$ where $w_{1}=y^{-1}x^{-1}yxy$.
Theorem 6.
For each fixed $n>0$, the 1-parameter family of presentations $\text{MS}(n,w_{k})$ are all AC-equivalent. Here, $w_{k}=y^{-k}x^{-1}yxy$ is parameterized by $k\in\mathbb{Z}$.
Proof.
Starting with the presentation,
$\text{MS}(n,w_{k})=\langle x,y\mid x^{-1}y^{n}x=y^{n+1},x=y^{-k}x^{-1}yxy\rangle,$
we will show that for any fixed $n$ and $k$, $\text{MS}(n,w_{k})$ is AC-equivalent to $\text{MS}(n,w_{k+1})$.
First, note that the first relation in $\text{MS}(n,w_{k})$ can be rearranged in the following three ways:
  1. (i)
    Multiplying by $y^{k-n}x$ from the left and by $y^{-1}$ from the right gives
    $y^{k}xy^{-1}=y^{k-n}xy^{n}.$
  2. (ii)
    Multiplying by $y^{-1}$ from the left and by $x^{-1}y^{-1}$ from the right, and inverting the relation gives
    $y^{-(n-1)}xy=yxy^{-n}.$
  3. (iii)
    Multiplying by $y^{-1}x$ from the left and by $y^{k-n}$ from the right, and inverting the relation gives
    $y^{n-k}x^{-1}y^{-(n-1)}=y^{-(k+1)}x^{-1}y.$
Now, we can rearrange the second relation to get $y^{k}xy^{-1}=x^{-1}yx$. Substituting (i) gives $y^{k-n}xy^{n}=x^{-1}yx$, which we can rewrite as $yxy^{-n}=xy^{k-n}x$. Now substituting (ii) gives $y^{-(n-1)}xy=xy^{k-n}x$, which we may rewrite as $y^{n-k}x^{-1}y^{-(n-1)}=xy^{-1}x^{-1}$. Finally, substituting (iii) gives $y^{-(k+1)}x^{-1}y=xy^{-1}x^{-1}$, which is equivalent to $x=y^{-(k+1)}x^{-1}yxy=w_{k+1}$. ∎
Theorem 7.
For each $n>0$ and $k\in\mathbb{Z}$, $\text{MS}(n,w_{k})$ is AC-equivalent to the presentation
$P(n,k)=\langle x,y\mid y^{n-k-1}x^{-1}yx=xyx^{-1}y^{n-k}\ ,\ x=y^{-k}x^{-1}yxy\rangle,$
of total length $|k|+|n-k|+|n-k-1|+11$. For each fixed $n$, this length is minimum when $k=n-1$, which gives the presentation
$P(n,n-1)=\langle x,y\mid x^{-1}yx=xyx^{-1}y\ ,\ x=y^{-(n-1)}x^{-1}yxy\rangle,$
of length $n+11$.
Proof.
Starting with the presentation,
$\text{MS}(n,w_{k})=\langle x,y\mid x^{-1}y^{n}x=y^{n+1},x=y^{-k}x^{-1}yxy\rangle,$
the second relation may be equivalently expressed as
  1. (i)
    $y^{k}xy^{-1}=x^{-1}yx$
  2. (ii)
    $y^{-1}xy^{k}=xyx^{-1}$
Multiplying the first relation by $y^{-1}x$ from the left and by $y^{-1}$ from the right, we get $y^{n-1}xy^{-1}=y^{-1}xy^{n}$, which we may rewrite as $y^{n-k-1}\left(y^{k}xy^{-1}\right)=\left(y^{-1}xy^{k}\right)y^{n-k}$. Substituting from (i) on the LHS and from (ii) on the RHS gives the required result. ∎
Combining the previous three results shows that, for each fixed $n$, $AK(n)$ is AC-equivalent to all members of the infinite families $\text{MS}(n,w_{k})$ and $P(n,k)$. From this, Theorem 4 follows immediately from setting $k=n-1$.
Remark 8.
While $P(n,n-1)$ is always the shortest presentation from the $P$ family, other $P(n,k)$ may also give reductions in length of $\text{AK}(n)$. To find which presentations of $P$-family are shorter than $\text{AK}(n)$, we compare the length of $\text{AK}(n)$, i.e. $72n+$ 7, with the length of $P(n,k)$ given by:
$\begin{cases}2n+10-k,&\text{if }k\leq n-1,\\ 3k+12-2n,&\text{if }k\geq n.\end{cases}$
This is less than $72n+$ 7 when $14\leq k\leq n-$ 1 or $n\leq k<\frac{1}{3}\left(4n-5\right)$.
We also note that the AC-triviality of $\text{AK}(2)$, when combined with Theorem 6, gives the following result.
Corollary 9.
Each member of the infinite families $\text{MS}(2,w_{k})$ and $P(2,k)$ is AC-trivial.
3.6. Limitations and extensions
@@
Another interesting direction to pursue involves changing the sorting function used by greedy search. Consider the family of greedy search algorithms where the priority queue is sorted by a fixed function $f:(r_{1},r_{2})\to\mathbb{R}$ which takes a presentation and outputs a real number [^8] We can also reinterpret breadth-first search in this manner if, given as input to $f$ the path length $l$, we have the function return $l$. From this perspective it becomes clear that, whenever breadth-first search succeeds, it finds a path of minimal length.. The greedy search algorithm used here is based on the function $f(r_{1},r_{2})\to|r_{1}|+|r_{2}|$. An interesting option would be to use machine learning to try and learn a good function $f$. This idea is inspired by search-based reinforcement learning systems, such as AlphaGo and AlphaZero, which combine Monte Carlo Tree Search with learned heuristics to evaluate and prioritize states dynamically. While these systems go beyond classical greedy search by incorporating exploration and planning through tree-based rollouts, the core principle of leveraging a learned heuristic function to guide search is analogous.
@@ 4. Reinforcement learning
This section is organized as follows: in Section 4.1, we discuss how the problem underlying the Andrews–Curtis conjecture can be formulated as a Markov Decision Process. In Section 4.2, we discuss details of a specific reinforcement learning algorithm, called the Proximal Policy Optimization algorithm, which we used to find AC trivializations of balanced presentations. Finally, in Section 4.3, we discuss the results of our work, comparing the performance of PPO with that of the classical search algorithms studied in the previous section.
@@ 4.1. Markov decision process
Figure 4. The basic RL cycle.
A schematic picture of how these objects interact with each other is as follows. We start with a state $s_{0}$ sampled from the distribution $\rho$ and take an action $a_{0}$. This results in a state $s_{1}$ with probability $P(s_{1}\mid s_{0},a_{0})$. The transition gets a “reward” $r_{0}=R(s_{0},a_{0},s_{1})$ which quantifies the effectiveness of the action in contributing toward achieving an ultimate goal. From state $s_{1}$, we repeat this process, obtaining a trajectory of states
@@
For a given problem at hand, we may not a priori know the actions $\{a_{t}\}$ and states $\{s_{t+1}\}$ that maximize the return. Deep reinforcement learning presents a solution to this problem: we train a neural network that learns a map from states to actions with the objective of maximizing the cumulative return. More precisely, we learn a map called the “policy” function $\pi\colon S\to\mathcal{P}(A)$ that assigns to each state a probability distribution over actions. At time step $t$, an action $a_{t}\sim\pi(\cdot\mid s_{t})$ is sampled that gives the next state $s_{t+1}$.999 In general, we need to specify the probability transition function $P(\cdot\mid s_{t},a_{t})$ from which the next state would be sampled. In this paper, we take this probability distribution to be a delta function centered at a single state, which we write as $a_{t}(s_{t})$. In the next subsection we discuss the specific algorithm and the objective function that we used in our study.
@@ 4.2. Proximal policy optimization
It turns out that this update depends on the gradient of the logarithm of the policy function itself and an “advantage function” $A^{\pi}(s,a)$ which quantifies the relative benefit of taking action $a$ in state $s$ under policy $\pi$ [SpinningUp2018].101010 Mathematically, the advantage function is the difference between “on-policy action-value function” $Q^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)\mid s_{0}=s,a_{0}=a]$ and the “on-policy value function”, $V^{\pi}(s)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)\mid s_{0}=s]$. These functions give an estimate of the cumulative return given that the current state-action pair is $(s,a)$ in the case of $Q^{\pi}$ and the current state is $s$ in the case of $V^{\pi}$. Explicitly,
@@
Here $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\text{old}}(a_{t}|s_{t})}$ represents the probability ratio, and $\epsilon$ is a small positive constant (usually set around 0.1 or 0.2) that controls the clipping value. The raw and the clipped ratio ensure that excessively large policy updates are curtailed, making the optimization process more stable.
The advantage function is estimated during the optimization process using the Generalized Advantage Estimation method of [schulman2018highdimensional]. This method requires an estimation of the on-policy value function (c.f. Footnote 10). In PPO, the actor-critic framework is used for this purpose. We train two neural networks in parallel, where the “actor” learns the policy $\pi_{\theta}$ and the “critic” learns the value function.111111 Sometimes, these neural networks are made to share their parameters, which helps in stability during training. We experimented with shared as well as unshared parameters and observed better performance by keeping the parameters independent. The value loss $L^{V}$ is the mean squared error between the values for a state as estimated by the critic network before and after a gradient update. Lastly, PPO adds entropy $S$ of the action-space distribution to the full loss function to prevent premature convergence to a suboptimal policy. The full loss function is as follows,
@@
In policy gradient algorithms such as PPO, we alternate between collecting data through the current policy in the “rollout phase” and updating the policy weights through gradient descent in the “optimization phase”. There are various hyperparameters associated to these phases, which we briefly describe now. The rollout phase involves generating multiple trajectories of the underlying Markov Decision Process in parallel using the current policy function. The hyperparameters of this phase include the number of parallel actors $N$, the rollout length $T^{\prime}$, i.e. the number of steps each actor takes, the discount factor $\gamma$ (c.f. Section 4.1) and a bootstrapping hyperparameter of Generalized Advantage Estimation $\lambda_{\text{GAE}}$. The dataset of $N\times T^{\prime}$ examples collected in this phase is then used to update the policy. In each epoch, the dataset is shuffled and split into mini-batches before performing gradient descent. In addition to the optimizer hyperparameters (such as learning rate schedule), the number of epochs and the size of mini-batches are hyperparameters of the optimization process. Further details of hyperparameters are summarized in Appendix A.
4.3. Application to the MS series
The set of all balanced presentations of the trivial group with a fixed number of generators and the set of AC-transformations play the roles of sets $S$ and $A$, respectively. (Recall notations introduced in Section 4.1). Once we choose a reward function $R$ and an initial state distribution $\rho$, we may use the Proximal Policy Optimization algorithm to learn the policy function $\pi$. We tested a few different reward functions in our experiments, observing that the following candidate led to the best performance and stability in training.